OpenStudy (anonymous):

3(x + 1) = -2(x - 1) + 6. Help please . I've got 3x+3=-2x-2+6

OpenStudy (johnweldon1993):

That's almost correct 3(x + 1) = -2(x - 1) + 6 3x + 3 yes..... - -2x + 2 + 6 Because -2 times -1 = 2 So we now have 3x + 3 = -2x + 2 + 6 Combine all the like terms...3x + 2x = 5x 5x + 3 = 2 + 6 2 + 6 = 8 so 5x + 3 = 8 subtract 3 from both sides... 5x = 5 divide both sides by 5 x = 1